if square root 1 xsquare X sign theta= x, then prove that Tan square theta + cot square theta = x square+1/xsquare

Grade:12th pass

1 Answers

Shashank Singh
6 years ago
v(1+x^2)sin theta = x
sin^2theta= x^2/(1+x^2)
and cos^2 theta= 1/(1+x^2)
so, tan^2 theta = x^2
cot^2 theta = 1/x^2
therefore, tan^2 theta + cot^2 theta = x^2 + 1/x^2............... [ans]